Cashews sell for $5 per pund and peanuts sell for $2 per pound. How many pounds of each would you use to make 25 pounds of a mixture that sells for $92?
Answer by checkley77(12844)

5x+2(25-x)=92
5x+50-2x=92
3x=92-50
3x=42
x=42/3
x=14 pounds of Cashews is used.
25-14=11 pounds of peanuts are used.
Proof:
5*14+2*11=92
70+22=92
92=92